# QZO: A Catalog of 5 Million Quasars from the Zwicky Transient Facility

### v1

@@ 6 Conclusion
We publish a catalog of reliable QSO candidates based on time domain observations, created using deep learning classification of ZTF light curves with a transformer model, and then ensembled by XGB with WISE survey data. The high quality sample consists of 4,849,574 QSOs, with an estimated QSO F1 score of 97%. For 33% of QZO objects, with available WISE data, we publish redshifts with estimated error $\Delta z/(1+z)=0.14$. We show the superiority of light curve based classification over static griz magnitudes and colors, and we reach similar classification capabilities to the WISE and Gaia surveys. However, both WISE and Gaia gain a significant improvement when optical light curves are added into their classifications. We also show that the ZTF light curve classifications are the most important features for QSO detection. In the case of red point-like objects detectable by both the Gaia and WISE surveys, $W1-W2$ is the most important feature, but the addition of light curve classification also improves the classification. We find that given at least 50 observation epochs, the QSO F1 scores deteriorate by 10pp between the objects at $5\sigma$ SNR magnitude limit $g\sim 20.8$ and the faintest data points at $g\sim 21.5$. We robustly classify objects fainter than the magnitude limit by requiring $g<\mathrm{n_{obs}}/80+20.375$. At the ZTF 3 day median cadence, a $\geq 900$ days time span is required to obtain a QSO F1 score higher than 90%. However, at the highest available time span of 1800 days, even a 12 day median cadence can provide similar results. Additionally, even a 500 day time span or 30 day median cadence can provide results higher than 80%.
@@
SJN and MJG are supported by the US National Science Foundation (NSF) through grant AST-2108402. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grants No. AST-1440341 and AST-2034437 and a collaboration including current partners Caltech, IPAC, the Oskar Klein Center at Stockholm University, the University of Maryland, University of California, Berkeley , the University of Wisconsin at Milwaukee, University of Warwick, Ruhr University Bochum, Cornell University, Northwestern University and Drexel University. Operations are conducted by COO, IPAC, and UW. Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, and NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology. WISE and NEOWISE are funded by the National Aeronautics and Space Administration. European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. Based on observations obtained at the Hale Telescope, Palomar Observatory, as part of a collaborative agreement between the Caltech Optical Observatories and the Jet Propulsion Laboratory. The Gordon and Betty Moore Foundation, through both the Data-Driven Investigator Program and a dedicated grant, provided critical funding for SkyPortal. All the PS data used in this paper can be found in MAST: https://doi.org/10.17909/55e7-5x63 (catalog 10.17909/55e7-5x63). All the AllWISE (Wright et al., 2019) and Gaia EDR3 (Collaboration, 2020) data used in this paper can be found in IPAC.
@@ Appendix A Published files
We publish the QZO catalog in the QZO.csv file, which includes 4,849,574 objects and columns as described in table 6, excluding the duplicate objects flag. The classifications are based on XGB models trained on ZTF g-band median magnitude and light curves classification with transformer model, as well as WISE W[1-4] magnitudes and colors. The photo-zs are based on ZTF g-band magnitude and WISE magnitudes and colors. We remove duplicated ZTF light curves by removing objects which within the full ZTF catalog have at least one neighbor within $1\arcsec$ with more ZTF observation epochs. The final number of quasars was achieved with magnitude, number of observation epochs, and minimum quasar classification probability cuts, such that $g<\mathrm{n_{obs}}/80+20.375$, and $p_{\mathrm{QSO}}>0.9$, where $p_{\mathrm{QSO}}$ is XGB classification probability for the quasar class. The photo-zs are available for 35% of these objects, depending on the availability of WISE observations.
In the file ZTF_all_QSO.csv, we publish the same columns from table 6, for 78,078,450 objects classified as QSOs by at least one of the two XGB models with and without WISE features. There are no cuts applied, and there are no duplicates removed. 26% of objects are marked with the duplicates flag.
We publish the train data predictions in the file train.csv. The file contains 2,588,221 records, with ZTF ID and duplicates flag missing. Picking the longest light curve for each SDSS object removes ZTF duplicates.
@@ Appendix B Misclassified light curves
Figures 16, 18 and 20 show ZTF g-band light curves of QSOs, galaxies and stars, respectively, classified as QSOs by the XGB model based on the transformer classifications and g-band median magnitude. This qualitative analysis is based on a small sample of light curves, but shows interesting results. Stars classified as quasars are variable but not periodic, while galaxies classified as quasars can be actually SDSS misclassified AGNs. The QSOs misclassified as galaxies and stars generally show low amplitude variability.

### v2

@@ 6 Conclusion
We publish a catalog of reliable QSO candidates based on time domain observations, created using deep learning classification of ZTF light curves with a transformer model, and then ensembled by XGB with WISE survey data. The high quality sample consists of 4,849,574 QSOs, with an estimated QSO F1 score of 97%. For 33% of QZO objects, with available WISE data, we publish redshifts with estimated error $\Delta z/(1+z)=0.14$. We show the superiority of light curve based classification over static griz magnitudes and colors, and we reach similar classification capabilities to the WISE and Gaia surveys. However, both WISE and Gaia gain a significant improvement when optical light curves are added into their classifications. We also show that the ZTF light curve classifications are the most important features for QSO detection. In the case of red point-like objects detectable by both the Gaia and WISE surveys, W1W2W1-W2 is the most important feature, but the addition of light curve classification also improves the classification. We find that given at least 50 observation epochs, the QSO F1 scores deteriorate by 10pp between the objects at $5\sigma$ SNR magnitude limit $g\sim 20.8$ and the faintest data points at $g\sim 21.5$. We robustly classify objects fainter than the magnitude limit by requiring $g<n_{\mathrm{obs}}/80+20.375$. At the ZTF 3 day median cadence, a $\geq 900$ days time span is required to obtain a QSO F1 score higher than 90%. However, at the highest available time span of 1800 days, even a 12 day median cadence can provide similar results. Additionally, even a 500 day time span or 30 day median cadence can provide results higher than 80%.
@@
Data availability
We publish the catalog and models on Zenodo (Nakoneczny et al., 2025), while the code is available on Zenodo (Nakoneczny, 2025), and GitHub  [^6] https://github.com/snakoneczny/ztf-agn.
@@
SJN and MJG are supported by the US National Science Foundation (NSF) through grant AST-2108402. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grants No. AST-1440341 and AST-2034437 and a collaboration including current partners Caltech, IPAC, the Oskar Klein Center at Stockholm University, the University of Maryland, University of California, Berkeley , the University of Wisconsin at Milwaukee, University of Warwick, Ruhr University Bochum, Cornell University, Northwestern University and Drexel University. Operations are conducted by COO, IPAC, and UW. Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, and NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology. WISE and NEOWISE are funded by the National Aeronautics and Space Administration. European Space Agency (ESA) mission Gaia (cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. Based on observations obtained at the Hale Telescope, Palomar Observatory, as part of a collaborative agreement between the Caltech Optical Observatories and the Jet Propulsion Laboratory. The Gordon and Betty Moore Foundation, through both the Data-Driven Investigator Program and a dedicated grant, provided critical funding for SkyPortal. PS data can be found in MAST (STScI, 2022), AllWISE (Wright et al., 2019) and Gaia EDR3 (Collaboration, 2020) data can be found in IPAC.
@@ Appendix A Published files
The QZO.csv file includes 4,849,574 objects, and provides columns as described in table 6, excluding the duplicate objects flag. The classifications are based on XGB models trained on ZTF g-band median magnitude and light curves classification with transformer model, as well as WISE W[1-4] magnitudes and colors. The photo-zs are based on ZTF g-band magnitude and WISE magnitudes and colors. We remove duplicated ZTF light curves by removing objects which within the full ZTF catalog have at least one neighbor within $1\arcsec$ with more ZTF observation epochs. The final number of quasars was achieved with magnitude, number of observation epochs, and minimum quasar classification probability cuts, such that $g<n_{\mathrm{obs}}/80+20.375$, and $p_{\mathrm{QSO}}>0.9$, where $p_{\mathrm{QSO}}$ is XGB classification probability for the quasar class. The photo-zs are available for 35% of these objects, depending on the availability of WISE observations.
File ZTF_all_QSO.csv provides all the columns for 78,078,450 objects classified as QSOs by at least one of the two XGB models with and without the WISE features. There are no cuts applied, and there are no duplicates removed. 26% of objects are marked with the duplicates flag.
We publish the train data predictions in the file train.csv. The file contains 2,588,221 records, with ZTF ID and duplicates flag missing. Selecting the longest ZTF light curve for each non duplicated SDSS object removed ZTF duplicates.
@@ Appendix B Misclassified light curves
Figures 21, 23 and 25 show ZTF g-band light curves of QSOs, galaxies and stars, respectively, classified as QSOs by the XGB model based on the transformer classifications and g-band median magnitude. This qualitative analysis is based on a small sample of light curves, but shows interesting results. Stars classified as quasars are variable but not periodic, while galaxies classified as quasars can be actually SDSS misclassified AGNs. The QSOs misclassified as galaxies and stars generally show low amplitude variability.